Q-Factor at resonance is given by ICA O ILLO VICO VLIV O
Answer #1

\text{ Ans}\text{ (d)}\text{ }\text{V}_\text{L}/\text{V}

Explanation:

We know the Q factor in LCR circuit is defined as the sharpness of resonance and it is defined as the ratio of the inductive reactance or capacitative reactance at resonance and resresistance.

\text{ Q}\text{ factor}=\frac{\omega_0 L}{R} = \frac{1}{\omega_0CR}

\text{or},\text{ Q}\text{ factor}=\frac{X_L}{R} = \frac{X_C}{R}

\text{or},\text{ Q}\text{ factor}=\frac{IX_L}{IR} = \frac{IX_C}{IR}=\frac{V_L}{V}=\frac{V_C}{V}

​​​​\text{so}\text{ we}\text{ get},\text{Q}\text{ factor}=\frac{V_L}{V}

why other options are not correct:

(a)

I_C=\frac{V}{X_C}=\frac{V}{\frac{1}{\omega_0 C}}=V\omega_0C\text{Q}\text{ factor}=\frac{1}{\omega_0 CR}=\frac{V}{\omega_0 CRV}=\frac{V}{I_cR}=\frac{I}{I_C}

So, Q=Ic/I cannot be corrected.

(b)

I_L=\frac{V}{X_L}=\frac{V}{\omega_0 L}

Q \text{ factor}=\frac{\omega_0L}{R}=\frac{V}{I_L R}=\frac{I}{I_L}

So, Q= Il/I cannot be corrected.

(c) from above calculation we see ,

Q=Vc/V

So Q=V/Vc cannot be corrected.